OpenStudy (anonymous):

If cos(x-3pi/2)+sin(3pi/2 + x) + sin(32pi + x) -18cos(19pi-x) + cos(56pi +x) -9sin(x+17pi) can be represented in the form of asinx + bcosx , then find (a+b)

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

Question updated

ganeshie8 (ganeshie8):

\(\large \cos(2n\pi +x) = \cos(x)\) \(\large \sin(2n\pi +x) = \sin(x)\)

OpenStudy (anonymous):

yes how we use that heere

OpenStudy (anonymous):

the first two terms we can use , first of all how to convert that into asinx + bcosx

OpenStudy (anonymous):

i thinking , it would be difficult on OS to do this sum, what u say better i ask my prof.

OpenStudy (anonymous):

i think**

OpenStudy (anonymous):

i don't know

ganeshie8 (ganeshie8):

cos(x-3pi/2)+sin(3pi/2 + x) + sin(32pi + x) -18cos(19pi-x) + cos(56pi +x) -9sin(x+17pi) simplifies to below after using the previous listed properties cos(x-3pi/2)+sin(3pi/2 + x) + sin(x) -18cos(pi-x) + cos(x) -9sin(x+pi)

OpenStudy (anonymous):

I will skip this question for now

ganeshie8 (ganeshie8):

why do you want to skip this question ? its so basic and your prof wont be impressed if you to to him with these silly questions